OpenStudy (anonymous):

there are 81 coins in a parking meter. There are 15 fewer quarters than dimes in a parking meter. The value of the coins is $13.05. How many dimes are there

OpenStudy (anonymous):

do you need help with this ?

OpenStudy (anonymous):

d + q = 81 --> q = 81 - d .10d + .25q = 13.05 now sub 81 - d in for q in the 2nd equation and solve for d